OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. 8 + 16 + 24 + . . . + 8n = 4n(n + 1)

OpenStudy (anonymous):

@bibby

OpenStudy (perl):

there are two parts to induction proof basis case and inductive step

OpenStudy (anonymous):

How do I solve this then?

OpenStudy (anonymous):

or like how do I answer this ?

OpenStudy (bibby):

assume the base case is true and prove it for k+1

OpenStudy (perl):

lets not skip the 'basis' case , (i mean)

OpenStudy (perl):

is this statement true for n=1 ?

OpenStudy (bibby):

what does that entail? prove it for n=1 or 2 and then prove it holds true for n+1?

OpenStudy (anonymous):

firstly put n=1 .and find out if its true or not

OpenStudy (perl):

prove it for n=1 , then prove that if true for k , then its true for k+1

OpenStudy (perl):

separately

OpenStudy (bibby):

I see. my bad

OpenStudy (loser66):

guys, finish your helps, please.

OpenStudy (perl):

it might be easier if we rewrite the expression using P(n)

OpenStudy (loser66):

@perl go to my post to help me, I can handle here

OpenStudy (perl):

check that 8*1 = 4(1)(1+1)

OpenStudy (bibby):

the series is the sum of n multiples of 8 so n=1: 8 n=2: 8+16 etc. prove that 8=(1)(4)(1+1)

OpenStudy (loser66):

basis step: past hypothesis step: assume it is true for k , that is 8+16+.....+8k = 4k(k+1) now, need prove it is true for k+1, that is

OpenStudy (loser66):

8+16+......+8k + 8(k+1)= 4(k+1)(k+1+1)=4(k+1)(k+2)

OpenStudy (perl):

hint: add 8(k+1) to both sides of that equation

OpenStudy (perl):

loser66, be careful about your right side of equation (that is what we want to prove)

OpenStudy (loser66):

oh, I think you are at my post @perl

OpenStudy (loser66):

I know, I just let it there to know what we have to get from the left hand side.

OpenStudy (loser66):